OpenStudy (anonymous):

If you came across an article which said scientists had discovered an organism that has three of each chromosome instead of two, would you be skeptical? Why or why not. I'm honestly so confused with this question.

OpenStudy (anonymous):

I'll medal if someone helps I'm super confused ://

OpenStudy (anonymous):

@Hannahfitz I would make no judgements based only on that statement. Do you have any other background info for us?

OpenStudy (anonymous):

@Hannahfitz However, I think I can help you with what the spirit of the question is interested in. Have you heard about copy number of chromosomes? The ploidy of the cell? Humans have a total of 46 chromosomes. Of those chromosomes, 44 are somatic chromosomes (non-sex chromosomes) and 2 are sex chromosomes (X and Y). Are we okay at this point? Okay, so although we have 46 chromosomes, we have pairs. So we have 23 pairs of chromosomes. Having pairs of chromosomes means we are diploid. This is frequently written as 2n. You can think of "n" has the number of pairs of chromosomes. So, literally \[2*n=2*23=46\]A total of 46 chromosomes. It is important to note that the pairs are why you hear about dominant and recessive genes. Colorblindness is a good example for this concept. Back to our pairs: there is one time when humans don't have a true pair, males XY. Female humans are XX and human males are XY. For all of our chromsomes, and other diploid organisms, a "pair" is like the female XX in humans. Now back to colorblindness- The gene(s) that deal with color vision are on the X chromosome and versions of the gene that do not work correctly are recessive. This means one copy is enough to have normal vision. For females, they have two X chromosomes and would have to have the "bad" version of the gene on both copies. Males only have one X and so we have a much high chance of being colorblind. We tend to think of most things as diploid like us, but the fact is there are plenty of organisms without paired chromosomes and some only have 1 chromosome. There are also organisms that don't have pairs, but instead have 4, 5, or even 18 copies of each chromosome. If you want to really blow your mind there are some fungi (in Dikarya) that have two or more distinct nuclei per cell. Sometimes one nucleus is 1n and the other is 2n or 3n. It is crazy. Hope that is more helpful than confusing.